The time married men with children spend on child care averages 9.8 hours per week (Time, March 12, 2012). You belong to a professional group on family practices that would like to do its own study to determine if the time married men in your area spend on child care per week differs from the reported mean of 9.8 hours per week. The standard deviation for the amount of time married men in your area spend on child care is believed to be a constant 3.9 hours per week. A sample of 200 couples produced a mean of 9.1 hours spents on child care per week.

If the true population mean is 9.8 hours, you made a

A. type i error B. type ii error c. correct decision
Answer #1

Test statistic

z= \frac{9.1-9.8}{3.9/\sqrt{200}}

z = -2.54

p - Value = 2*P(z < -2.54) = 2*0.00555 = 0.0111

This value is less than 0.05 so we reject Ho. Hence,

We are rejecting Ho when it is true. So, we are making Type I Error.